# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ 3.1. Properties of the weak IMCF
Let us collect in this section few known properties on the weak IMCF constructed in Theorem 3.1.
Let $(M,g)$ be a smooth complete Riemaniann manifold. Let $o\in M,R>0$ and let $w$ be given by Theorem 3.1. Let $T\in\mathbb{R}$ be such that $\{w\leq T\}\subset B_{R}(o)$ (such a $T$ exists by Theorem 3.1). For every $t\in(-\infty,T]$, set $E_{t}:=\{w<t\}$, and $E_{t}^{+}=\{w\leq t\}$. Then the following hold.
@@
    If $n\leq 7$, then for every $t\in(-\infty,T]$ both the set $E_{t}$ and the set $E_{t}^{+}$ have $C^{1,\alpha}$ boundary, for some $\alpha>0$. Indeed, as a direct consequence of Section 3, the previous sets are local $(\Lambda,r_{0})$-minimizers in $B_{R}(o)\setminus\{o\}$, see, e.g., [MaggiBook, Example 21.2]; thus by (the Riemannian analogue of) [MaggiBook, Theorem 21.8] one gets the sought claim. In fact, one can obtain $C^{1,1}$ regularity of the sublevel sets [Heidusch].
    It follows that $\partial E_{t}=\partial\{w\leq t\}=\{w=t\}$ for almost every $t\leq T$, and that $\partial E_{s}\to\partial E_{t}$ in $C^{1}$ as $s\nearrow t$, for every $t\leq T$, cf. [HuiskenIlmanen, Eq. (1.10)].
    Moreover, it is meaningful to speak about the weak mean curvature $H$ on $\partial E_{t}$, see, e.g., [HuiskenIlmanen, page 16]. Moreover, $H=|\nabla w|>0$ $\mathcal{H}^{n-1}$-almost everywhere on $\partial E_{t}$ for almost every $t\in(-\infty,T]$, see [HuiskenIlmanen, Equation (1.12)], and [HuiskenIlmanen, Lemma 5.1].
@@
    For almost every $t\in(-\infty,T]$, the boundary $\partial E_{t}$ has weak second fundamental form $A$ (see [GerochMonotonicityMoser, Definition 1.3], or [HuiskenIlmanen, pages 401–405]), and
@@
    The last inequality is a consequence of [GerochMonotonicityMoser, Theorem 1.1(vi)].
@@
[…]
Let $(M,g)$ be a smooth complete Riemaniann manifold of dimension $n=3$. Let $o\in M,R>0$ and let $w$ be given by Theorem 3.1. Denote $E_{t}\vcentcolon=\{w<t\}$. Let $T$ be such that $\{w\leq T\}\subset B_{R}(o)$. Let $H$ be the weak mean curvature of the boundary $\partial E_{t}$. Then the following hold.
@@
    Then $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to-\infty$. Moreover, for every $-\infty<r<s\leq T$, if $\partial E_{t}$ is connected for every $t\in[r,s]$, then
@@
Denote $E_{t}^{+}\vcentcolon=\{w\leq t\}$ and let $E^{p}_{t}\vcentcolon=\{w^{2R}_{p}\leq t\}$ for $w^{2R}_{p}$ as in Theorem 3.1. We shall omit the superscript $2R$ in the sequel, as $R$ is fixed. We know from Theorem 3.1 that $w_{p}$ converges to $w$ locally uniformly on $B_{2R}(o)\setminus\{o\}$ along a sequence $p_{i}\to 1$.
We claim that for any $\sigma\in(-\infty,T),\varepsilon>0$ there exists $\tilde{p}>1$ such that $E^{p_{i}}_{t-\varepsilon}\subset E_{t}^{+}\subset E^{p_{i}}_{t+\varepsilon}$ for any $p_{i}\in(1,\tilde{p})$ and $t\in[\sigma,T]$.
We prove the containment $E^{p_{i}}_{t-\varepsilon}\subset E_{t}^{+}$ first. Assume by contradiction that there exist $\sigma\in(-\infty,T)$, $\varepsilon>0$ such that, up to subsequence, there exist $t_{i}\in[\sigma,T]$ such that $E^{p_{i}}_{t_{i}-\varepsilon}\not\subset E_{t_{i}}^{+}$ for any $i$. Then there exist points $x_{i}\in B_{2R}(o)\setminus\{o\}$ such that $w_{p_{i}}(x_{i})\leq t_{i}-\varepsilon$ but $w(x_{i})>t_{i}$ for any $i$. Hence $\liminf_{i}\mathsf{d}(x_{i},o)>0$, for otherwise $-\infty=\liminf_{i}w(x_{i})\geq\sigma$ by Theorem 3.1. Moreover, there exists $\eta>0$ such that $w_{p_{i}}>T+\eta$ on $\partial B_{R}(o)$ for large $i$, for otherwise $\{w\leq T\}\cap\partial B_{R}(o)$ would be nonempty. Hence $x_{i}\in B_{R}(o)$ for large $i$, for if $x_{i}\in B_{2R}(o)\setminus\overline{B}_{R}(o)$, since $w_{p_{i}}(x_{i})\leq T-\varepsilon$ and $w_{p_{i}}(x)\to+\infty$ as $r(x)\to 2R$, then $w_{p_{i}}$ would have an interior minimum on $B_{2R}(o)\setminus\overline{B}_{R}(o)$, contradicting the maximum principle for $p$-harmonic functions. Hence, up to subsequence, $t_{i}\to\tau\in[\sigma,T]$ and $x_{i}\to x\in\overline{B}_{R}(o)\setminus\{o\}$. But this contradicts the local uniform convergence, as this implies $\tau-\varepsilon\geq\lim_{i}w_{p_{i}}(x_{i})=w(x)=\lim_{i}w(x_{i})\geq\tau$.
@@
Recalling Section 3, for any $\sigma\in(-\infty,T),\varepsilon>0$ there exists $\tilde{p}>1$ such that
@@
for any $p_{i}\in(1,\tilde{p})$ and $t\in[\sigma,T]$. Letting first $p_{i}\to 1$ and then $\varepsilon\to 0$ and $\sigma\to-\infty$, we find that
@@
Fix $t\in(-\infty,T]$. It follows from the very definition of weak IMCF as in [HuiskenIlmanen, Minimizing Hull Property 1.4] that $E_{t}^{+}$ is strictly outward minimizing relatively to $B_{2R}(o)$; meaning that whenever $E_{t}^{+}\subsetneq F\subset\subset B_{2R}(o)$, then $P(E_{t}^{+})<P(F)$. It is readily checked that the proof of [FogagnoloMazzieriHulls, Theorem 1.2] can be localized in the ball $B_{2R}(o)$, thus yielding that
@@
Finally, as in [HuiskenIlmanen, Minimizing Hull Property 1.4(iv)], there holds $P(E_{t}^{+})=P(E_{t})$, so that (3.27) follows.
We show the second property in (3.24). By applying (3.23) with $\overline{r}=2r(x)$ and sending $p=p_{i}\to 1^{+}$, we get that there exists $\tilde{\eta},\vartheta>0$ such that if $r(x)<\tilde{\eta}$, then $w(x)\leq 2\log(r(x))+\vartheta$. On the other hand, for $\bar{t}$ small enough we have $E_{t}\subset B_{e^{(t+C)/2}}(o)$ for every $t<\bar{t}$, see item (4) in Theorem 3.1. Up to possibly taking a smaller $\bar{t}$, for every $t<\bar{t}$ we have $\partial E_{t}\subset\{w=t\}\subset M\setminus B_{e^{(t-\vartheta)/2}}$. Thus
@@
The fact that $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to-\infty$ is a direct consequence of the first item and the definition of Hawking mass. The last part of the second item is a consequence of the analogue of [HuiskenIlmanen, Geroch Monotonicity Formula 5.8] in our setting. A detailed proof the Geroch Monotonicity Formula for weak IMCFs defined through limits of $p$-harmonic functions as in our setting will be provided in the forthcoming [BenattiPludaPozzetta]. ∎
@@ 3.2. Connectedness of level sets of the weak IMCF
In the following lemma we collect some facts about connectedness of level set of the weak IMCF. This is analogous to [HuiskenIlmanen, Connectedness Lemma 4.2], we provide a proof for the convenience of the reader.
[…]
Let $\Omega$ be an open set with Lipschitz boundary in a smooth complete $n$-dimensional Riemannian manifold $(M,g)$ with $n\leq 7$, and let $o\in\Omega$. Let $\Omega\subset\subset U$, where $U$ is an open set, and let $w\in\mathrm{Lip}_{\mathrm{loc}}(U\setminus\{o\})$ be a weak solution of the IMCF on $U\setminus\{o\}$, see Section 3. Then the following hold.
@@
$M$$U$$\Omega$$o$$\Omega_{t_{1}}$$\Omega_{t_{0}}$$\Omega^{\prime\prime}$
Figure 1. The picture sketches the assumptions of Section 3.2.
Proof of Section 3.2.
The proof of item (1) follows verbatim as in [HuiskenIlmanen, Connectedness Lemma 4.2(i)]. Let us repeat it here for the ease of the reader. Assume by contradiction a connected component $C$ of $\{w>t\}\cap\Omega$ is relatively compact in $\Omega\setminus\{o\}$. Then consider the function $v:=w$ on $U\setminus(C\cup\{o\})$ and $v:=t$ on $C$. By (3.1) we get
@@
Now assume by contradiction a connected component $C^{\prime}$ of $\{w<t\}\cap\Omega$ is relatively compact in $\Omega\setminus\{o\}$. Take $\bar{t}:=\min_{\overline{C^{\prime}}}w$. Then for $10<\eta<$ 1 small enough there is a connected component $C^{\prime\prime}$ of $\{w<\bar{t}+\eta\}\cap\Omega$ inside $C^{\prime}$. Notice that $w\geq\bar{t}>\bar{t}+\eta-1$ on $C^{\prime\prime}$. Repeating the previous argument with $\bar{t}+\eta$ in place of $t$ and $C^{\prime\prime}$ in place of $C$ gives again a contradiction.
@@
The item (3) is inspired by [HuiskenIlmanen, Connectedness Lemma 4.2(ii)]. Analogous arguments have appeared in [MunteanuWang, Lemma 2.3], [JDGLiTam, Lemma 6.1], and [LeeBook, Lemma 4.46]. We give here a self-contained proof using item (2) and the Mayer–Vietoris sequence.
Recall from item (1) of Section 3.1 that $\partial\{w<t\}$ is $C^{1,\alpha}$ for every $t\in(t_{0},t_{1})$, and that $\partial\{w<t\}=\{w=t\}$ for almost every $t\in(t_{0},t_{1})$. It is sufficient to prove that $\partial\{w<t\}$ is connected for $t$ such that $\partial\{w<t\}=\{w=t\}$. Indeed, for $\tau\in(t_{0},t_{1})$ such that $\partial\{w<\tau\}\neq\{w=\tau\}$, there exists a sequence $t_{i}\nearrow\tau$ such that $\partial\{w<t_{i}\}=\{w=t_{i}\}$. From item (1) of Section 3.1 we know that $\partial\{w<t_{i}\}\to\partial\{w<\tau\}$ in $C^{1}$, hence connectedness will be preserved in the limit.
Hence we can assume by contradiction that there exists $t\in(t_{0},t_{1})$ such that $\partial\{w<t\}=\{w=t\}$ is not connected. Since $\partial\{w<t\}$ is $C^{1,\alpha}$, it has a finite number $m\geq 2$ of connected components. Since $\bigcap_{\eta>0}\{t-\eta<w<t+\eta\}=\{w=t\}=\partial\{w<t\}$, there exists $\eta$ small enough such that $[t-\eta,t+\eta]\subset(t_{0},t_{1})$ and $\{t-\eta<w<t+\eta\}$ has $m^{\prime}\geq 2$ connected components. Call $A:=\{w<t+\eta\}\cup\{o\}$ and $B:=\{w>t-\eta\}\cap\Omega$. Notice that $A$ and $B$ are open, and by item (2) they are both connected.
@@ 4. Proof of the main results
In this section we prove the main theorems Theorem 1.3, Theorem 1.1.
@@ 4.1. Producing a set satisfying the reverse Euclidean isoperimetric inequality
In this section we show how, in the hypotheses of Theorem 1.1 and Theorem 1.2, we can produce sets that satisfy the Euclidean reverse isoperimetric inequality with sharp constant.
[…]
Let $(M,g)$ be a smooth complete Riemannian manifold of dimension $3$. Let $o\in M$ and $R>0$. Let $w$ be given by Theorem 3.1. Let $T\in\mathbb{R}$ be such that $\{w<T\}\subset\subset B_{\rho}(o)$ for some $\rho\leq R$. Suppose that $\partial\{w<t\}$ is connected for any $t<T$. If $R_{g}\geq-\delta$ on $B_{\rho}(o)$, for some $\delta>0$, then
@@
We recall that $E_{t}$ has $C^{1,\alpha}$ boundary, $\partial E_{t}$ admits weak mean curvature $H$ for all $t<T$, and $H=|\nabla w|>0$ $\mathcal{H}^{2}$-a.e. for a.e. $t\in(-\infty,T)$, see item (1) of Section 3.1. The Hawking mass of a set $\Omega$ with $C^{1}$ boundary $\partial\Omega$ possessing weak mean curvature $H$ is given by
@@
As $\{w\leq T^{\prime}\}\subset\subset B_{\rho}(o)$ for any $T^{\prime}<T$, we can apply Section 3.1, which yields that $P(E_{t})=4\pi e^{t}$ for any $t<T$, and that $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to-\infty$. Hence the Geroch Monotonicity formula (3.26) with $r\to-\infty$ and $s=t$, together with the fact that $R_{g}\geq-\delta$ on $B_{\rho}(o)$, gives that for all $t\in(-\infty,T)$ there holds
@@
Since $\mathfrak{m}_{H}(\partial E_{t})=\frac{2\pi^{1/2}e^{t/2}}{64\pi^{3/2}}\left(4\pi-\int_{\partial E_{t}}\frac{H^{2}}{4}\right)$, for every $t\in(-\infty,T)$, we find
@@
$\begin{split}|E_{t}|&\geq\int_{E_{t}\cap\{|\nabla w|>0\}}\frac{1}{|\nabla w|}|\nabla w|=\int_{-\infty}^{t}\left(\int_{\partial E_{\tau}\cap\{|\nabla w|>0\}}\frac{1}{|\nabla w|}\right)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{t}\left(\int_{\partial\Omega_{\tau}}\frac{1}{|\nabla w|}\right)\,\mathrm{d}\tau\geq\int_{-\infty}^{t}\left(P(\Omega_{\tau})^{3/2}\left(\int_{\partial\Omega_{\tau}}|\nabla w|^{2}\right)^{-1/2}\right)\\ &\geq\int_{-\infty}^{t}\frac{2\pi e^{3t/2}}{\sqrt{1+\frac{2}{3}\delta e^{t}}}\geq\int_{-\infty}^{t}\frac{2\pi e^{3t/2}}{\sqrt{1+\frac{2}{3}\delta e^{T}}}=\frac{1}{\sqrt{1+\frac{2}{3}\delta e^{T}}}\,\frac{4}{3}\pi e^{3t/2}=\frac{1}{\sqrt{1+\frac{2}{3}\delta e^{T}}}\,\frac{P(E_{t})^{3/2}}{6\sqrt{\pi}}.\end{split}$ (4.4)
In the $C^{0}_{\rm loc}$-asymptotically flat case we construct sets with arbitrarily large perimeter and volume that satisfy the reverse Euclidean isoperimetric inequality. This represents the crucial step for the proof of Theorem 1.1 and of Theorem 1.3.
[…]
Let $(M,g)$ be a $3$-dimensional complete $C^{0}$-Riemannian manifold without boundary, let $K\subset M$ be a compact set, and let $\Omega$ be an unbounded connected component of $M\setminus K$. Assume that $\Omega$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$ (Section 1.1), and that $R_{g}\geq 0$ in the approximate sense on $\Omega\setminus K^{\prime}$ (Section 1.1), where $K^{\prime}\subset M$ is a compact set. Then, there exists a universal constant $\vartheta\in(0,1)$ such that the following holds.
@@
Let $\rho>1$, $\eta<1/2$ to be chosen. The choice of $\rho$, only depending on $\mathscr{P}$ and on geometric constants on $\mathbb{R}^{3}$, will be made clear during the proof in (4.10). We do not insist on the precise choice of $\eta$ for the sake of readability; however, it will be clear from the proof that choosing $\eta<10^{-3}$ is sufficient. In this proof we will repeatedly use the elementary metric results recorded in Section 2.1, and in the last part of the proof of Section 2.3.
By applying a contradiction argument and Section 2.1 (see also the beginning of the proof of Section 2.3) we can find a compact set $\mathcal{C}\supset(K\cup K^{\prime})$ such that for every $x\in\Omega\setminus\mathcal{C}$ there exists $F_{x}:(B_{64\rho+64}(x),g)\to(\mathbb{R}^{3},g_{\mathrm{eu}})$ which is a $(1+\eta)$-biLipschitz diffeomorphism with its image, with $F_{x}(x)=0$, and whose image contains $B_{32\rho+32}^{\mathbb{R}^{3}}(0)$. Let us now fix $o$ such that
@@
Notice that (see Section 2.1) the compact $\mathcal{C}$ can be chosen such that, additionally, we have

### v3

@@ 3.1. Properties of the weak IMCF
Let us collect in this section few known properties on the weak IMCF constructed in Theorem˜3.6.
Let $(M,g)$ be a smooth complete Riemaniann manifold. Let $o\in M,R>0$ and let $w$ be given by Theorem˜3.6. Let $T\in\mathbb{R}$ be such that $\{w\leq T\}\subset B_{R}(o)$ (such a $T$ exists by Theorem˜3.6). For every $t\in(-\infty,T]$, set $E_{t}:=\{w<t\}$, and $E_{t}^{+}=\{w\leq t\}$. Then the following holds.
@@
    If $n\leq 7$, then for every $t\in(-\infty,T]$ both the set $E_{t}$ and the set $E_{t}^{+}$ have $C^{1,\alpha}$ boundary, for some $\alpha>0$. Indeed, as a direct consequence of Section˜3, the previous sets are local $(\Lambda,r_{0})$-minimizers in $B_{R}(o)\setminus\{o\}$, see, e.g., [52, Example 21.2]; thus by (the Riemannian analogue of) [52, Theorem 21.8] one gets the sought claim. In fact, one can obtain $C^{1,1}$ regularity of the sublevel sets [35].
    It follows that $\partial E_{t}=\partial\{w\leq t\}=\{w=t\}$ for almost every $t\leq T$, and that $\partial E_{s}\to\partial E_{t}$ in $C^{1}$ as $s\nearrow t$, for every $t\leq T$, cf. [38, Eq. (1.10)].
    Moreover, it is meaningful to speak about the weak mean curvature $H$ on $\partial E_{t}$, see, e.g., [38, page 16]. Moreover, $H=|\nabla w|>0$ $\mathcal{H}^{n-1}$-almost everywhere on $\partial E_{t}$ for almost every $t\in(-\infty,T]$, see [38, Equation (1.12)], and [38, Lemma 5.1].
@@
    For almost every $t\in(-\infty,T]$, the boundary $\partial E_{t}$ has weak second fundamental form AA (see [57, Definition 1.3], or [38, pages 401–405]), and
@@
    The last inequality is a consequence of [57, Theorem 1.1(vi)].
@@
Proposition 3.14.
Let $(M,g)$ be a smooth complete Riemaniann manifold of dimension $n=3$. Let $o\in M,R>0$ and let $w$ be given by Theorem˜3.6. Denote $E_{t}\vcentcolon=\{w<t\}$. Let $T$ be such that $\{w\leq T\}\subset B_{R}(o)$. Let $H$ be the weak mean curvature of the boundary $\partial E_{t}$. Then the following holds.
@@
    Then $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to$-$\infty$. Moreover, for a.e. $\infty<r<s\leq T$, if $\partial E_{t}$ is connected for a.e. $t\in[r,s]$, then
@@
Denote $E_{t}^{+}\vcentcolon=\{w\leq t\}$ and let $E^{p}_{t}\vcentcolon=\{w^{2R}_{p}\leq t\}$ for $w^{2R}_{p}$ as in Theorem˜3.6. We shall omit the superscript $2R$ in the sequel, as $R$ is fixed. We know from Theorem˜3.6 that $w_{p}$ converges to $w$ locally uniformly on $B_{2R}(o)\setminus\{o\}$ along a sequence $p_{i}\to 1$.
[…]
@@
Recalling Section˜3, for any $\sigma\in(-\infty,T),\varepsilon>0$ there exists $\tilde{p}>1$ such that
@@
[…]
@@
Fix $t\in(-\infty,T]$. It follows from the very definition of weak IMCF as in [38, Minimizing Hull Property 1.4] that $E_{t}^{+}$ is strictly outward minimizing relatively to $B_{2R}(o)$; meaning that whenever $E_{t}^{+}\subsetneq F\subset\subset B_{2R}(o)$, then $P(E_{t}^{+})<P(F)$. It is readily checked that the proof of [30, Theorem 1.2] can be localized in the ball $B_{2R}(o)$, thus yielding that
@@
Finally, as in [38, Minimizing Hull Property 1.4(iv)], there holds $P(E_{t}^{+})=P(E_{t})$, so that (3.27) follows.
We show the second property in (3.24). By applying (3.23) with $\overline{r}=2r(x)$ and sending $p=p_{i}\to 1^{+}$, we get that there exists $\tilde{\eta},\vartheta>0$ such that if $r(x)<\tilde{\eta}$, then $w(x)\leq 2\log(r(x))+\vartheta$. On the other hand, for $\bar{t}$ small enough we have $E_{t}\subset B_{e^{(t+C)/2}}(o)$ for every $t<\bar{t}$, see item (4) in Theorem˜3.6. Up to possibly taking a smaller $\bar{t}$, for every $t<\bar{t}$ we have $\partial E_{t}\subset\{w=t\}\subset M\setminus B_{e^{(t-\vartheta)/2}}$. Thus
@@
The fact that $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to$-$\infty$ is a direct consequence of the first item and the definition of Hawking mass.
The last part of the second item is the analogue of [38, Geroch Monotonicity Formula 5.8], which provides the analogous monotonicity formula for a weak IMCF globally defined on a complete manifold. The validity of (3.26) in the present setting is well-known to the experts, and can be deduced as a consequence of [14, Theorem 4.1]. In fact, the estimate (3.26), which does not takes into account all the positive terms present in the complete [38, Geroch Monotonicity Formula 5.8] (see also [14, Theorem 5.5]), can also be derived by suitably passing to the limit $p\to 1$ in [9, Theorem 2.11], taking into account [14, Corollary 4.7, Corollary 4.8]. ∎
@@ 3.2. Connectedness of level sets of the weak IMCF
In the following lemma we collect some facts about connectedness of level set of the weak IMCF. This is analogous to [38, Connectedness Lemma 4.2], we provide a proof for the convenience of the reader.
Lemma 3.15.
Let $\Omega$ be an open set with Lipschitz boundary in a smooth complete $n$-dimensional Riemannian manifold $(M,g)$ with $n\leq 7$, and let $o\in\Omega$. Let $\Omega\subset\subset U$, where $U$ is an open set, and let $w\in\mathrm{Lip}_{\mathrm{loc}}(U\setminus\{o\})$ be a weak solution of the IMCF on $U\setminus\{o\}$, see Section˜3. Then the following hold.
@@
[…]
Figure 1. The picture sketches the assumptions of Section˜3.2.
Proof of Section˜3.2.
The proof of item (1) follows verbatim as in [38, Connectedness Lemma 4.2(i)]. Let us repeat it here for the ease of the reader. Assume by contradiction a connected component $C$ of $\{w>t\}\cap\Omega$ is relatively compact in $\Omega\setminus\{o\}$. Then consider the function $v:=w$ on $U\setminus(C\cup\{o\})$ and $v:=t$ on $C$. By (3.1) we get
@@
Now assume by contradiction a connected component $C^{\prime}$ of $\{w<t\}\cap\Omega$ is relatively compact in $\Omega\setminus\{o\}$. Take $\bar{t}:=\min_{\overline{C^{\prime}}}w$. Then for $0<\eta<1$ small enough there is a connected component $C^{\prime\prime}$ of $\{w<\bar{t}+\eta\}\cap\Omega$ inside $C^{\prime}$. Notice that $w\geq\bar{t}>\bar{t}+\eta-1$ on $C^{\prime\prime}$. Repeating the previous argument with $\bar{t}+\eta$ in place of $t$ and $C^{\prime\prime}$ in place of $C$ gives again a contradiction.
@@
The item (3) is inspired by [38, Connectedness Lemma 4.2(ii)]. Analogous arguments have appeared in [58, Lemma 2.3], [51, Lemma 6.1], and [49, Lemma 4.46]. We give here a self-contained proof using item (2) and the Mayer–Vietoris sequence.
Recall from item (1) of Section˜3.1 that $\partial\{w<t\}$ is $C^{1,\alpha}$ for every $t\in(t_{0},t_{1})$, and that $\partial\{w<t\}=\{w=t\}$ for almost every $t\in(t_{0},t_{1})$. It is sufficient to prove that $\partial\{w<t\}$ is connected for $t$ such that $\partial\{w<t\}=\{w=t\}$. Indeed, for $\tau\in(t_{0},t_{1})$ such that $\partial\{w<\tau\}\neq\{w=\tau\}$, there exists a sequence $t_{i}\nearrow\tau$ such that $\partial\{w<t_{i}\}=\{w=t_{i}\}$. From item (1) of Section˜3.1 we know that $\partial\{w<t_{i}\}\to\partial\{w<\tau\}$ in $C^{1}$, hence connectedness will be preserved in the limit.
Hence we can assume by contradiction that there exists $t\in(t_{0},t_{1})$ such that $\partial\{w<t\}=\{w=t\}$ is not connected. Since $\partial\{w<t\}$ is $C^{1,\alpha}$, it has a finite number $m\geq 2$ of connected components. Since $\bigcap_{\eta>0}\{t-\eta<w<t+\eta\}=\{w=t\}=\partial\{w<t\}$, there exists $\eta$ small enough such that $t-\eta,t+\eta]\subset(t_{0},t_{1})$ and $\{t-\eta<w<t+\eta\}$ has $m^{\prime}\geq 2$ connected components. Call $A:=\{w<t+\eta\}\cup\{o\}$ and $B:=\{w>t-\eta\}\cap\Omega$. Notice that $A$ and $B$ are open, and by item (2) they are both connected.
@@ 4. Proof of the main results
In this section we prove the main theorems Theorem˜1.6, Theorem˜1.4.
@@ 4.1. Producing a set satisfying the reverse Euclidean isoperimetric inequality
In this section we show how, in the hypotheses of Theorem˜1.4 and Theorem˜1.5, we can produce sets that satisfy the Euclidean reverse isoperimetric inequality with sharp constant.
Lemma 4.1.
Let $(M,g)$ be a smooth complete Riemannian manifold of dimension $3$. Let $o\in M$ and $R>0$. Let $w$ be given by Theorem˜3.6. Let $T\in\mathbb{R}$ be such that $\{w<T\}\subset\subset B_{\rho}(o)$ for some $\rho\leq R$. Suppose that $\partial\{w<t\}$ is connected for any $t<T$. If $R_{g}\geq-\delta$ on $B_{\rho}(o)$, for some $\delta>0$, then
@@
We recall that $E_{t}$ has $C^{1,\alpha}$ boundary, $\partial E_{t}$ admits weak mean curvature $H$ for all $t<T$, and $H=|\nabla w|>0$ $\mathcal{H}^{2}$-a.e. for a.e. $t\in(-\infty,T)$, see item (1) of Section˜3.1. The Hawking mass of a set $\Omega$ with $C^{1}$ boundary $\partial\Omega$ possessing weak mean curvature $H$ is given by
@@
As $\{w\leq T^{\prime}\}\subset\subset B_{\rho}(o)$ for any $T^{\prime}<T$, we can apply Section˜3.1, which yields that $P(E_{t})=4\pi e^{t}$ for any $t<T$, and that $\mathfrak{m}_{H}(\partial E_{t})\to 0$ for almost every $t\to$-$\infty$. Hence the Geroch Monotonicity formula (3.26) along a sequence $r\to$-$\infty$ and $s=t$, together with the fact that $R_{g}\geq-\delta$ on $B_{\rho}(o)$, gives that for a.e. $t\in(-\infty,T)$ there holds
@@
Since $\mathfrak{m}_{H}(\partial E_{t})=\frac{2\pi^{1/2}e^{t/2}}{64\pi^{3/2}}\left(4\pi-\int_{\partial E_{t}}\frac{H^{2}}{4}\right)$, for a.e. $t\in(-\infty,T)$, we find
@@
$\begin{split}|E_{t}|&\geq\int_{E_{t}\cap\{|\nabla w|>0\}}\frac{1}{|\nabla w|}|\nabla w|=\int_{-\infty}^{t}\left(\int_{\partial E_{\tau}\cap\{|\nabla w|>0\}}\frac{1}{|\nabla w|}\right)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{t}\left(\int_{\partial E_{\tau}}\frac{1}{|\nabla w|}\right)\,\mathrm{d}\tau\geq\int_{-\infty}^{t}\left(P(E_{\tau})^{3/2}\left(\int_{\partial E_{\tau}}|\nabla w|^{2}\right)^{-1/2}\,\mathrm{d}\tau\right)\\ &\geq\int_{-\infty}^{t}\frac{2\pi e^{3\tau/2}}{\sqrt{1+\frac{2}{3}\delta e^{\tau}}}\,\mathrm{d}\tau\geq\int_{-\infty}^{t}\frac{2\pi e^{3\tau/2}}{\sqrt{1+\frac{2}{3}\delta e^{T}}}\,\mathrm{d}\tau\\ &=\frac{1}{\sqrt{1+\frac{2}{3}\delta e^{T}}}\,\frac{4}{3}\pi e^{3t/2}=\frac{1}{\sqrt{1+\frac{2}{3}\delta e^{T}}}\,\frac{P(E_{t})^{3/2}}{6\sqrt{\pi}}.\end{split}$ (4.4)
In the $C^{0}_{\rm loc}$-asymptotically flat case we construct sets with arbitrarily large perimeter and volume that satisfy the reverse Euclidean isoperimetric inequality. This represents the crucial step for the proof of Theorem˜1.4 and of Theorem˜1.6.
Proposition 4.2.
Let $(M,g)$ be a $3$-dimensional complete $C^{0}$-Riemannian manifold without boundary, let $K\subset M$ be a compact set, and let $\Omega$ be an unbounded connected component of $M\setminus K$. Assume that $\Omega$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$ (Section˜1.1), and that $R_{g}\geq 0$ in the approximate sense on $\Omega\setminus K^{\prime}$ (Section˜1.1), where $K^{\prime}\subset M$ is a compact set. Then, there exists a universal constant $\vartheta\in(0,1)$ such that the following holds.
@@
Let $\rho>1$, $\eta<1/2$ to be chosen. The choice of $\rho$, only depending on $\mathscr{P}$ and on geometric constants on $\mathbb{R}^{3}$, will be made clear during the proof in (4.10). We do not insist on the precise choice of $\eta$ for the sake of readability; however, it will be clear from the proof that choosing $\eta<10^{-3}$ is sufficient. In this proof we will repeatedly use the elementary metric results recorded in Section˜2.1, and in the last part of the proof of Section˜2.3.
By applying a contradiction argument and Section˜2.1 (see also the beginning of the proof of Section˜2.3) we can find a compact set $\mathcal{C}\supset(K\cup K^{\prime})$ such that for every $x\in\Omega\setminus\mathcal{C}$ there exists $F_{x}:(B_{64\rho+64}(x),g)\to(\mathbb{R}^{3},g_{\mathrm{eu}})$ which is a $(1+\eta)$-biLipschitz diffeomorphism with its image, with $F_{x}(x)=0$, and whose image contains $B_{32\rho+32}^{\mathbb{R}^{3}}(0)$. Let us now fix $o$ such that
@@
Notice that (see Section˜2.1) the compact $\mathcal{C}$ can be chosen such that, additionally, we have